OpenStudy (anonymous):

2(3x + 5) = 46 could somebody answer that for me? please? i don't understand these at all ;-;

OpenStudy (anonymous):

46/2 = 23 so 3x+5 has to equal 23 23-5=18 3 times what = 18 ?

OpenStudy (anonymous):

6..so the answers 6?

OpenStudy (anonymous):

plug in 6 for x and see

OpenStudy (anonymous):

don't insult me :c please...im stupid. and don't know what to do..if i put 6 as x it would be 36....at least thats what i understand

OpenStudy (anonymous):

3*6=18 18+5=23 23*2=46

OpenStudy (anonymous):

-facepalm- sorry..i get it now..thank you. you helped a lot

OpenStudy (anonymous):

haha no problem